OpenStudy (anonymous):

could anybody please explain how to find pedal equation of of : r=a(e^thetacot(α))

OpenStudy (anonymous):

Do you know what pedal equation is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If you explain what it is it might help to figure the question out.

OpenStudy (anonymous):

do u know about tangents and normals in polar coordinates

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

it is a equation of curve in terams of r and p only where p =rsinΦ and Φ is angle b/w the radius vector of a point and the tangent at that point

OpenStudy (anonymous):

"terams" ---- terms

OpenStudy (anonymous):

The equation you wrote is ambiguous.

OpenStudy (anonymous):

Is p not \(\rho\) from spherical coordinates?

OpenStudy (anonymous):

Strange notation nonetheless.

OpenStudy (anonymous):

i am posting a photograph of an example in the book

OpenStudy (anonymous):

\[ r=a(e^\theta \cot(\alpha)) \]Is this where we are starting?

OpenStudy (anonymous):

The fact that we have \(\alpha\) is a bit annoying.

OpenStudy (anonymous):

Is the \(\cot\) supposed to be in the exponent?

OpenStudy (anonymous):

yes cot is in exponent and α is constant

OpenStudy (anonymous):

First thing you do is differentiate \(r\) with respect to \(\theta\)

OpenStudy (anonymous):

\[ \frac{dr}{d\theta }=a\cot(\alpha)e^{\theta \cot(\alpha)} \]

OpenStudy (anonymous):

\[ \tan(\phi)=\frac{ae^{\theta \cot\alpha }}{a\cot\alpha e^{\theta \cot\alpha }}=\frac{1}{\cot\alpha }=\tan \alpha \]

OpenStudy (anonymous):

So I guess \(\phi =\alpha\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

tell further

OpenStudy (anonymous):

\[ \ln(r)-\ln(a) = \theta \cot\alpha \]

OpenStudy (anonymous):

\[ \cot \alpha = \frac{\cos \alpha}{\sin\alpha } = \frac{\sqrt{1-\sin^2\alpha }}{\sin \alpha }= \frac{\sqrt{1-\sin^2\phi}}{\sin \phi} \]

OpenStudy (anonymous):

\[ r\cot \alpha =\frac{\sqrt{1-P^2}}{P} \]

OpenStudy (anonymous):

\[ r\ln r-r\ln a = \theta \frac{\sqrt{1-P^2}}{P} \]

OpenStudy (anonymous):

That theta doesn't seem to be going anywhere.

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @wio \[ r\cot \alpha =\frac{\sqrt{1-P^2}}{P} \] \(\color{blue}{\text{End of Quote}}\) This should be \[ r\cot \alpha =\frac{\sqrt{r^2-P^2}}{P} \]